### v1

# Bulk plasmons in elemental metals

Abstract
The spectral properties, momentum dispersion, and broadening of bulk plasmonic excitations of 25 elemental metals are studied from first principles calculations in the random-phase approximation. Spectral band structures are constructed from the resulting momentum- and frequency-dependent inverse dielectric function. We develop an effective analytical representation of the main collective excitations in the dielectric response, by extending our earlier model based on multipole-Padé approximants (MPA) to incorporate both momentum and frequency dependence. With this tool, we identify plasmonic quasiparticle dispersions exhibiting complex features, including non-parabolic energy and intensity dispersions, discontinuities due to anisotropy, and overlapping effects that lead to band crossings and anti-crossings. We also find good agreement between computed results and available experiments in the optical limit. The results for elemental metals establish a reference point that can guide both fundamental studies and practical applications in plasmonics and spectroscopy.
I Introduction
Elemental metals, composed of a single type of metallic atom, are foundational to both modern technology and condensed matter physics [1, 2]. Examples include alkali metals (e.g., Na, K), alkaline earth metals (e.g., Mg, Ca), and transition or noble metals (e.g., Al, Cu, Ag, Au). These materials are characterized by high electrical and thermal conductivity, a strong but ductile metallic bonding, and a conduction band partially filled with delocalized electrons. Delocalized electrons respond collectively to external fields and are responsible for the characteristic optical and electronic properties that distinguish metals from insulators and semiconductors. Thus, understanding the electronic properties of elemental metals, including their collective excitations, is of significant theoretical and practical importance. Metals with a relatively simple electronic band structure, like Al and alkali metals, host delocalized conduction electrons that behave nearly as a free-electron gas [3, 4, 5, 6]. This property, along with a substantial amount of available experimental data, makes them ideal systems for studying fundamental electronic excitations and their underlying many-body interactions.
Collective oscillations of the electron density give rise to plasmonic quasiparticles. Their origin lies in the long-range Coulomb restoring force acting on charge-density fluctuations, which produces well-defined modes in the free-electron gas and, in real materials, dispersions modified by band structure and dielectric screening [7, 8, 9]. Plasmons can be excited by electromagnetic perturbations or incident charged particles, and appear as bulk or surface modes [10, 11]. Plasmons are central in several technological applications, including nanophotonics [12, 13], surface-enhanced spectroscopy [14], energy harvesting [15, 16], and emerging quantum plasmonic technologies [17, 18, 19]. Plasmons can also be engineered by modifying the dielectric environment or the electronic structure, for example through alloying  [20, 21].
Bulk plasmons correspond to longitudinal charge-density oscillations and are therefore optically inactive. They are observed in experiments such as electron energy-loss spectroscopy (EELS) [10, 11], where they appear as well-defined peaks in the measured spectra. Other techniques such as inelastic x-ray scattering and reflection electron energy loss spectroscopy (REELS) are also used to provide complementary insights into plasmonic properties [22, 23, 24, 25, 26, 27, 28]. In simple metals such as Al or Na, the plasmon peak is sharp and symmetric, akin to that of a free-electron gas [5, 10]. In contrast, in noble metals, such as Ag and Au, the occupied $d$-bands introduce inter-band transitions that broaden and shift the plasmon resonance, and give rise to much richer spectra [29, 4]. Such experimental signatures are broadly employed for material characterization, from determining free-carrier concentrations, phases, and thicknesses to exploring nanoplasmonic effects and probing the electronic band structure [27, 11].
Although the dielectric properties of elemental metals in the optical limit have been extensively studied and tabulated [29, 4, 5, 8, 28], properties at finite momentum are not as widely available. In this work, we report the frequency and momentum dependent inverse dielectric functions of 25 elemental metals computed from first principles within the random phase approximation (RPA) level of the many-body theory, using Kohn–Sham Bloch (KS) states obtained from density functional theory (DFT). We also provide an effective analytical representation of the main collective excitations in the inverse dielectric function, by generalizing our previous model based on multipole-Padé approximants (MPA) [30, 31, 32] through the inclusion of both momentum and frequency dependence. The generalized MPA model in two variables, MPA($\mathbf{q}$}), is used to describe the main spectral properties of such excitations.
The paper is organized as follows: Sec. II details the theoretical framework (Sec. II.1), the MPA and MPA($\mathbf{q}$}) representations (Sec. II.2), the set of studied materials (Sec. II.3), and the computational details of the first-principle calculations (Sec. II.4). Sec. III presents the results for the optical limit of intra-band contributions (Sec. III.1), the comparison between RPA and the experimental spectra (Sec. III.2), and the deviations from the free-electron gas model (Sec. III.3), while results at finite momentum are used for the construction of spectral band structures (Sec. III.4), and the generalized MPA($\mathbf{q}$}) model (Sec. III.5). Finally, Sec. IV holds the conclusions.
II Methods
II.1 Theoretical framework
The complex dielectric function $\varepsilon$ governs the response of a material to external electromagnetic fields. It depends on momentum $\mathbf{q}$} and frequency $\omega$. In the optical or long-wavelength limit, ${\mathbf{q}}\to 0$, the dielectric function of the free-electron gas reduces to the Drude form, characterized by a single pole [33, 4, 34, 5, 35]. Within the RPA approximation, Lindhard theory extends this classical Drude model to include the $\mathbf{q}$} dependence [9, 8], providing a first-order quantum-mechanical description of the free-electron response. The dielectric response in real materials, often deviates from the homogeneous fee-electron behavior, and is therefore much richer. The RPA dielectric function can be computed by first-principle calculations using KS-DFT states. In such an approach, the independent-particle microscopic polarizability is given by
$\chi_{0\mathbf{GG^{\prime}}}(\mathbf{q},\omega)=\sum_{n,m}\int_{\mathrm{BZ}}\frac{d\mathbf{k}}{(2\pi)^{3}}\rho^{*}_{nm\mathbf{k}}(\mathbf{q},\mathbf{G})\rho_{nm\mathbf{k}}(\mathbf{q},\mathbf{G^{\prime}})\times\\ f_{n\mathbf{k-q}}(1-f_{m\mathbf{k}})\left[\frac{2\Omega^{\mathrm{KS}}_{nm\mathbf{k}\mathbf{q}}}{\omega^{2}-(\Omega^{\mathrm{KS}}_{nm\mathbf{k}\mathbf{q}})^{2}}\right],$ (1)
where the $n$ and $m$ indices run over the bands, $\rho_{nm\mathbf{k}}(\mathbf{q},\mathbf{G})\equiv\langle n\mathbf{k}|e^{i(\mathbf{q}+\mathbf{G})\cdot\mathbf{r}}|m\mathbf{k-q}\rangle$ are transition matrix elements, the $f$ factors are the occupations of the KS states, $\Omega^{\mathrm{KS}}_{nm\mathbf{k}\mathbf{q}}=(\epsilon_{m\mathbf{k}}-\epsilon_{n\mathbf{k-q}})-i\delta$ are KS single-particle transitions, and the limit $\delta\to 0^{+}$ is implicit and ensures the correct time ordering [6].
The interacting microscopic polarizability and the inverse dielectric function are obtained by the Dyson equation:
$\chi$ $=\chi_{0}+\chi_{0}v\chi$ (2)
$\varepsilon^{-1}$ $=1+v\chi,$ (3)
where $v$ is the Coulomb potential. Even though the macroscopic polarizability, $\chi(\mathbf{q},\omega)$, is given by the $\mathbf{G}=\mathbf{G}^{\prime}=0$ component of the microscopic $\chi$, it is affected by the coupling of $\mathbf{G},\mathbf{G}^{\prime}\neq 0$ components through Eq. (2). These contributions are commonly referred to as local-field effects, as they account for microscopic variations of the induced fields within the crystal. Their inclusion accounts for the difference between the independent-particle approximation (IPA) and RPA.
In metallic systems, Eq. (1) encompasses intra- and inter-band transitions that mix in the Dyson equation of Eq. (2). We can thus separate both contributions to the interacting polarizability as
$\chi^{\mathrm{intra}}$ $=\chi-\chi^{\mathrm{inter}}$ (4)
$\chi^{\mathrm{inter}}$ $=\chi_{0}^{\mathrm{inter}}+\chi_{0}^{\mathrm{inter}}v\chi^{\mathrm{inter}}.$ (5)
This separation carries over to the dynamic part of the inverse dielectric function, $Y\equiv v\chi=\varepsilon^{-1}-1$, resulting in [31]
$Y({\mathbf{q}},\omega)=Y_{\mathrm{intra}}({\mathbf{q}},\omega)+Y_{\mathrm{inter}}({\mathbf{q}},\omega),$ (6)
where $Y_{\mathrm{intra}}\equiv v\chi^{\mathrm{intra}}$ and $Y_{\mathrm{inter}}\equiv v\chi^{\mathrm{inter}}$. Even if $\chi^{\mathrm{intra}}$ vanishes in the ${\mathbf{q}}\to 0$ limit, $Y_{\mathrm{intra}}$ remains finite due to the divergence of the Coulomb potential $v$. Therefore, the plasmonic quasiparticles observed in $Y({\mathbf{q}},\omega)$ at any value of $\mathbf{q}$} are a mixture of intra- and inter-band contributions.
An important property of the imaginary part of $Y$, or loss function, is the f-sum rule, which can be used to define the intra-band, inter-band, and plasma frequencies:
$\omega_{\mathrm{intra}}^{2}({\mathbf{q}})$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y_{\mathrm{intra}}({\mathbf{q}},\omega)]|$ (7)
$\omega_{\mathrm{inter}}^{2}({\mathbf{q}})$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y_{\mathrm{inter}}({\mathbf{q}},\omega)]|$ (8)
$\omega_{\mathrm{pl}}^{2}$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y({\mathbf{q}},\omega)]|.$ (9)
From Eq. (6) it follows that $\omega_{\mathrm{pl}}^{2}=\omega_{\mathrm{intra}}^{2}({\mathbf{q}})+\omega_{\mathrm{inter}}^{2}({\mathbf{q}})$. Whereas $\omega_{\mathrm{pl}}$ is $\mathbf{q}$}-independent, the respective intra- and inter-band contributions are $\mathbf{q}$} dependent, but with a dependence constrained by this condition.
The interacting $\chi$, and hence also $Y$, has a pole structure similar to that of $\chi_{0}$ in Eq. (1). However, the large number of single-particle transitions $\Omega^{\mathrm{KS}}_{vc\mathbf{k}\mathbf{q}}$ are mixed in the Dyson equation in Eq. (2), typically resulting in much simpler envelope functions [31]. For many systems, $Y$ in the optical limit ${\mathbf{q}}\to 0$ resembles the response of the free-electron gas [36, 37, 8], and can therefore be approximated by a single plasmon-pole model (PPA):
$Y^{\mathrm{PPA}}({\mathbf{q}}\to 0,\omega)=\frac{\Omega_{0}^{2}}{\omega^{2}-\Omega_{0}^{2}},$ (10)
where the pole is located at the plasmon energy, $\mathrm{Re}[\Omega_{0}]=\omega_{\mathrm{pl}}$. In the free-electron gas the plasmon energy is related to the electronic density, $\rho_{e}$, as (Gaussian units)
$\omega_{\mathrm{pl}}=\sqrt{4\pi\rho_{e}}.$ (11)
The extension of the PPA model to finite $\mathbf{q}$} is given by [38, 39, 40, 41, 42, 43, 44, 45]
$Y^{\mathrm{PPA}}({\mathbf{q}},\omega)=\frac{2R({\mathbf{q}})\Omega({\mathbf{q}})}{\omega^{2}-[\Omega({\mathbf{q}})]^{2}}\,.$ (12)
where the pole has a quadratic dispersion according to the Drude model, $\Omega(q)\approx\Omega_{0}(1+\Omega^{\prime\prime}q^{2}/2)$, with an effective electron mass factor $\Omega^{\prime\prime}$. The dispersion of the spectral weight is linked to the position of the pole by the f-sum rule of Eq. (9):
$\Omega_{0}^{2}=2R({\mathbf{q}})\Omega({\mathbf{q}}),$ (13)
which implies that $R({\mathbf{q}})$ decreases with the inverse of $\Omega({\mathbf{q}})$, as $\propto 1/q^{2}$.
II.2 Simple analytical representation: the multipole approach (MPA)
The RPA approximation is at the core of ab initio many-body theories beyond DFT, such as $GW$; however, expressing the operators in terms of single-particle excitations, as in the Lehmann representation, is both computationally impractical and makes difficult the physical interpretation of collective features. Alternatively, analytical models simplifying the frequency dependence of the dielectric function, starting from PPA, have been used to reduce the computational cost of first-principle calculations. However, the use of PPA is subject to accuracy limitations [30, 31]. More complex models, such as Padé approximants and many-pole schemes, have often been used to fit and interpret individual experimental spectra [46, 47, 48, 49, 50, 25, 51]. Within ab initio $GW$ calculations, effective MPA representations have been shown to provide an efficient description of the dielectric response, achieving an accuracy comparable to that of full-frequency approaches [30, 31, 32]. Such models generalize the PPA inverse dielectric response of Eq. (12), as follows
$Y^{\mathrm{MPA}}({\mathbf{q}},\omega)=\sum_{p}^{n_{Y}}\frac{2R_{p}({\mathbf{q}})\Omega_{p}({\mathbf{q}})}{\omega^{2}-[\Omega_{p}({\mathbf{q}})]^{2}},$ (14)
where the number of poles $n_{Y}$ is typically around 10.
In practice, poles and residues are obtained, for each $\mathbf{q}$}, by interpolation with a coarse sampling in the complex frequency plane [30, 31]. Despite considering each $\mathbf{q}$} separately, the procedure provides a good description of the $\mathbf{q}$} dispersion of the main poles [31]. MPA has also been combined with numerical methods to accelerate convergence with respect to the $\mathbf{q}$}-mesh, in the so-called constant approximation (CA) of Ref. [31] and the Monte Carlo $W$ averaging (W-av) of Refs. [52, 53].
In this work, we further advance this type of representation by proposing a generalized MPA model with an explicit $\mathbf{q}$} dependence, denoted MPA($\mathbf{q}$}). We consider a polynomial dispersion of the poles and residues, following the Taylor expansion around ${\mathbf{q}}=0$ of the independent-particle response and the free-electron gas. In practice, we truncate the expansion at the third order and consider specific $\mathbf{q}$}-directions:
$\begin{split}R_{p}(q)&\approx R_{p}(1+R^{\prime}_{p}q+R^{\prime\prime}_{p}q^{2}/2+R^{\prime\prime\prime}_{p}q^{3}/6)\\ \Omega_{p}(q)&\approx\Omega_{p}(1+\Omega^{\prime}_{p}q+\Omega^{\prime\prime}_{p}q^{2}/2+\Omega^{\prime\prime\prime}_{p}q^{3}/6),\end{split}$ (15)
where $R_{p}$ and $\Omega_{p}$ are the residues and pole positions in the optical limit $q\to 0$ along the selected direction, and the respective pairs $R^{\prime}_{p}$ and $\Omega^{\prime}_{p}$, $R^{\prime\prime}_{p}$ and $\Omega^{\prime\prime}_{p}$, and $R^{\prime\prime\prime}_{p}$ and $\Omega^{\prime\prime\prime}_{p}$ are the corresponding linear, quadratic, and cubic coefficients of the polynomial expansions. Although other $q$ dependencies can be explored, a third-order polynomial is shown to be sufficiently flexible to obtain accurate representations with relatively few poles, from 2 to 15, for all the systems studied here. Therefore, the generalized MPA($\mathbf{q}$}) model results in a quite compact representation of the momentum and frequency dependent $Y({\mathbf{q}},\omega)$, with the total number of parameters given by $8\times n_{Y}$. All the parameters are determined in a non-linear fit with physical constraints on the position of the poles (see Refs. [30, 31]):
$\begin{split}&0<\mathrm{Re}[\Omega_{p}({\mathbf{q}})]<\omega_{\mathrm{max}}\\ &-\mathrm{Re}[\Omega_{p}({\mathbf{q}})]<\mathrm{Im}[\Omega_{p}({\mathbf{q}})]<0.\end{split}$ (16)
Another relevant constraint is the f-sum rule, which for MPA($\mathbf{q}$}) takes the following form
$\omega_{\mathrm{pl}}^{2}=\sum_{p}^{n_{Y}}2\mathrm{Re}[R_{p}({\mathbf{q}})\Omega_{p}({\mathbf{q}})].$ (17)
However, this is strictly respected only when the integration interval over energy in Eq. (9) is sufficiently large. Therefore, we do not explicitly impose it for small energy ranges.
II.3 Reference data for the set of elemental metals
The dielectric properties of metals in the optical limit are widely studied and reported, e.g., in several handbooks and websites [54, 4]. However, the frequency dependent dielectric function is not always available and different sources can be inconsistent with respect to precision, energy range, etc. Experimental techniques such as EELS only measure the imaginary part of the response and need careful post-processing to retrieve the inverse dielectric function, which is nontrivial at finite momentum [55].
Here we consider reference experimental data from several sources. Data corresponding to the optical limit ${\mathbf{q}}\to 0$ were taken from the REELS measurements of Ref. [26] (set of 17 elemental metals: Ag, Au, Cu, Ni, Pb, Pt, Pd, Fe, Mo, Ta, V, W, Co, Ti, Zn, Bi, and Te), the EELS atlas [54] (for Be, Al, V, Cr, Cu, Zn, Ag, Sn, Te, W, Os, Pt, Au, Tl, Pb, Bi), and Ref. [56, 57] for Ca. In the case of EELS, we use a standard procedure, based on a fit to a power law model [58], to remove the background intensity from the zero-loss peak [11]. Reference data at finite momentum are available for Li [22, 23, 59], Be [24, 60], Na [61, 62], Mg [63], Al [63, 64, 65, 66, 67, 68, 62], and K [61], and for Cu, Ag, and Au [69].
A complete list of the 25 elemental metals studied here is shown in Table 1. The table includes the atomic number, the electronic configuration of valence electrons, the symmetry group, the experimental lattice parameters [70], and the volume of the unit cell. The computed intra- and inter-band frequencies are also provided.
II.4 Computational details
The crystalline structure of elemental metals is well characterized experimentally and theoretically [71]. In this work, we use the experimental geometries and lattice parameters listed in Table 1. DFT calculations were performed using the plane wave implementation of the Quantum Espresso package [72, 73] with the Perdew-Burke-Ernzerhof (PBE) variant of the GGA functional [74]. We adopted the norm-conserving optimized Vanderbilt pseudopotentials of Ref. [75], with a kinetic plane wave energy cutoff of 70 Ry. In the DFT calculations, the Brillouin zone was sampled with a $36\times 36\times 36$ Monkhorst-Pack grid for cubic, and $36\times 36\times 24$ for hexagonal materials. The calculations of dielectric spectra within RPA were performed with the yambo code [76, 77].
III Results
System $Z$ val. electrons space group (num.) lattice par. (Å) unit vol. (Å3) $\omega_{\mathrm{intra}}$ (eV) $\omega_{\mathrm{inter}}$ (eV) $\omega_{\mathrm{intra}}^{2}/\omega_{\mathrm{pl}}^{2}$ (%) $Z_{\mathrm{eff}}$
Li 3 2s1 Im_3m (229) 3.51 21.6218 6.56 4.66 66.5 0.72
Be 4 2s2 P63/mmc (194) 2.2858,   3.5843 8.1093 5.35 17.83 8.2 2.20
Na 11 3s1 Im_3m (229) 4.2906 39.4934 5.62 1.42 94.0 0.90
Mg 12 3s2 P63/mmc (194) 3.2094,  5.2108 23.2409 6.97 7.39 47.1 1.76
Al 13 3s2 3p1 fC Fm_3m (225) 4.0495 16.6014 11.84 9.22 62.2 2.63
K 19 4s1 Im_3m (229) 5.328 75.6245 4.15 10.58 13.4 0.76
Ca 20 4s2 Fm_3m (225) 5.5884 43.6317 4.20 15.08 7.2 2.10
Ti 22 3d2 4s2 P63/mmc (194) 2.9508,  4.6855 17.6659 3.74 26.48 2.0 3.77
V 23 3d3 4s2 Im_3m (229) 3.03 13.9091 7.96 29.61 6.7 4.22
Cr 24 3d5 4s1 Im_3m (229) 2.91 12.3211 6.80 32.72 4.1 4.85
Fe 26 3d6 4s2 Im_3m (229) 2.8665 11.7768 6.40 32.15 3.8 5.48
Co 27 3d7 4s2 P63/mmc (194) 2.5071,  4.0695 11.0761 6.00 31.90 3.4 5.43
Ni 28 3d8 4s2 Fm_3m (225) 3.524 10.9408 6.56 34.09 3.6 7.94
Cu 29 3d10 4s1 Fm_3m (225) 3.6149 11.8094 8.61 31.65 6.9 5.76
Zn 30 3d10 4s2 P63/mmc (194) 2.6649,  4.9468 15.2120 9.14 29.96 8.5 4.00
Mo 42 4d5 5s1 Im_3m (229) 3.147 15.5833 8.51 32.41 6.5 6.68
Pd 46 4d10 Fm_3m (225) 3.8907 14.7239 6.72 34.90 3.6 10.17
Ag 47 4d10 5s1 Fm_3m (225) 4.0853 17.0456 8.84 34.48 6.2 7.57
Sn 50 4d10 5s2 5p2 I41/amd (141) 5.8318,  3.1819 27.0540 8.85 29.45 8.3 3.35
Ta 73 4f14 5d2 6s2 Im_3m (229) 3.3013 17.9897 8.21 47.77 2.9 5.82
W 74 4f14 5d3 6s2 Im_3m (229) 3.1652 15.8553 6.91 50.89 1.81 6.85
Pt 78 4f14 5d9 6s1 Fm_3m (225) 3.9242 15.1075 8.38 38.66 4.5 12.05
Au 79 4f14 5d10 6s1 Fm_3m (225) 4.0782 16.9569 8.69 36.78 5.3 7.35
Tl 81 4f14 5d10 6s2 6p1 P63/mmc (194) 3.4566,  5.5248 28.5835 6.72 32.65 4.1 2.17
Pb 82 4f14 5d10 6s2 6p2 Fm_3m (225) 4.9508 30.3365 8.94 32.43 7.1 3.32
Table 1: List of properties of the 25 elemental metals studied in this work. The experimental lattice parameters and the corresponding unit cell volume are taken from the compilation made in Ref. [70].
III.1 Intra-band v.s. inter-band contributions
Fig. 1 shows the inter- and intra-band contributions to the plasma frequency as a function of the energy for Na, Ca, and Cu, computed at at the optical limit with Eqs. (7), (8) and (9). Inter-band contributions from semi-core and core states appear at large energies, e.g., above $25$ eV for the case of Na, and were not considered when evaluating the inter-band and plasma frequencies.
Figure 1: Spectral contributions of the intra-band (dashed orange line), the inter-band (dash-dotted blue line), and the plasma (solid green line) frequency as a function of energy, evaluated with Eqs. (7), (8), and (9), respectively. The total area under the plasma curve is filled with each corresponding contribution.
Table 1 reports the respective intra- and inter-band frequencies, and the percentage of intra-band contributions to the plasma frequency for the 25 studied metals. While the intra-band frequencies of all the metals range from 3.74 eV for Ti to $11.84$ eV for Al, the inter-band frequencies have a far wider range, from only $1.42$ eV for Na to $50.89$ eV for W. This huge span is due to inter-band contributions of d- and f-orbitals, thus intra-band contributions are relatively more relevant for metals with only $s$ and $p$ electrons. The plasmon of Na is dominated by intra-band contributions (94.0 %), followed by Li (66.5 %), and Al (62.2 %). Next in the list are Mg, with less than half (47.1 %), and K (13.4 %). The percentage for the rest of the metals is much lower.
III.2 Spectra in the optical limit
Figure 2: Comparison of the computed RPA loss function of V (a), Cu (b), and Zn (c) in the optical limit, with experimental EELS data from Ref. [54], REELS from Ref. [26] and theoretical IPA calculations from Ref. [26].
In Fig. 2 we compare the computed RPA loss function of V, Cu, and Zn in the optical limit, with EELS measurements from the EELS Atlas [54] and REELS data and IPA from Ref. [26]. These are examples of materials with complex screening effects resulting in several peaks in their response function, that require a description beyond a single pole PPA model [57, 4, 31].
The raw EELS data from Ref. [54] includes a background intensity from the zero-loss peak, which we removed using a standard power law model [58, 11], while the remaining intensity was normalized according to the maximum RPA intensity. These measurements have a lower resolution and a larger instrumental broadening than those of Ref. [26], however, for most cases, the energy range is larger, providing information on the tail of the spectra. The experimental loss function of Ref. [26] was obtained by performing Kramers-Kronig transformations [26, 11] on the raw REELS data. The results in the tail region are therefore affected by the finite energy interval in the frequency integral of such transformations, and in several cases, like Cu, are less accurate than the direct EELS measurements.
In all cases, both our RPA calculations and the IPA results from Ref. [26] are able to accurately describe the main features in the experiments. The inclusion of local field effects with RPA improves the description of the intensity with respect to IPA. Refining the DFT starting point, e.g., with DFT+U or hybrid functionals, may improve the description of the finer features in the loss function of materials with $d$-states and strong spin-orbit interactions [78, 79, 80, 55]. Such refinements are beyond the scope of this work.
Figure 3: Real (a) and imaginary (b) parts of the inverse dielectric function of V computed with RPA in the optical limit. The numerical data is compared with its corresponding MPA model with $n_{Y}=13$ poles. Dashed lines in the bottom panel represent the individual contributions of each pole. The most prominent pole is highlighted with a gray filling.
Fig. 3 compares the numerical RPA results for the real and imaginary parts of $\varepsilon^{-1}({\mathbf{q}}\to 0,\omega)$ of V in a smaller energy range, up to $30$ eV, with the fitted MPA model of Eq. (14) with $n_{Y}=13$. Even with such a small number of poles, the MPA model accurately reproduces the main features of the inverse dielectric response. The bottom panel shows dashed curves representing the individual contributions of each pole. Many of them carry a significant fraction of the total spectral weight and present a large broadening. The overlap of the poles illustrates the complex nature of plasmonic excitations in elemental metals like V.
III.3 Deviation from the free-electron gas and PPA
Figure 4: (a) Relation between the effective number of electrons, $Z_{\mathrm{eff}}$, and the number of valence electrons, $Z_{\mathrm{val}}$, listed in Table 1 for the set of studied metals. $Z_{\mathrm{eff}}$ is computed from the energy position of the main MPA pole, with Eq. (18). The black dashed line represents the identity relation. (b) Relation between the spectral weight, $\mathrm{Re}[2R]$, and the energy position, $\mathrm{Re}[\Omega]$, of the main MPA pole of all the studied metals. The black dashed line corresponding to $\mathrm{Re}[2R]=\mathrm{Re}[\Omega]$, represents the ideal PPA model where all the spectral weight is concentrated in a single pole. The presence of other poles with significant spectral weights lowers the position with respect to this PPA line. The gray dashed line represents a pole with half the spectral weight.
The free electron gas model has often been used to describe simple metals like Al and Na [1, 2, 81], considering an electronic density given by the number of their valence electrons. Such an approach can be used for other metals, but it is not always clear which electronic density to consider. With MPA in place, we can define an effective number of electrons from the expression of the classical plasmon energy in Eq. (11):
$Z_{\mathrm{eff}}\equiv\frac{\mathrm{Re}[\Omega]^{2}V}{4\pi},$ (18)
where $\Omega$ is the main pole of the MPA representation and $V$ is the volume of the unit cell.
Fig. 4(a) shows the relation between the computed $Z_{\mathrm{eff}}$ and the number of valence electrons, $Z_{\mathrm{val}}$, for the set of 25 studied metals. The values are listed in the last column of Table 1. For some of the cases $Z_{\mathrm{eff}}$ is close to the number of electrons in the outer orbitals. For example, the value of $0.9$, $2.1$, and $10.17$ are very close to $1$, 22, and $10$, corresponding to the number of electrons in the 3s1, 4s2, and 4d10 orbitals of Na, Ca, and Pd, respectively. However, in several cases like Co, Ag, and Ta, the mixed nature of the different orbital contributions to the plasmon energy and thus $Z_{\mathrm{eff}}$, prevents us from establishing any simple association.
On the ab initio side, the single pole PPA model, has been extensively used in $GW$ and similar approaches beyond DFT, even for systems with several prominent poles in their response functions. The MPA representation captures such a multipole behavior, and the spectral weight of its dominant pole provides a quantitative measure of how far is the PPA description.
Fig. 4 shows the spectral weight and the energy position of the main MPA pole in the response function of the 25 studied metals. According to Eq. (10), in PPA the spectral weight should equal the plasmon energy, represented by the black dashed line. The gray dashed line indicates half of the total spectral, $\mathrm{Re}[2R]=\mathrm{Re}[\Omega]/2$. The results show that Na, Mg, Al, K, and Sn are the systems well described by PPA, and the plasmon of Al has the largest spectral weight of all the metals. For Be, Li, Ca, W, Ta, Pb, and Tl the main pole has nearly half the total spectral weight, while for the remaining metals has even smaller weights. Ag has the smallest spectral weight and one of the largest deviations from PPA.
III.4 Spectral $Y({\mathbf{q}},\omega)$ band structures
Fig. 5 shows the dispersion of the main excitations in $Y({\mathbf{q}},\omega)$, along two high-symmetry $\mathbf{q}$}-lines of bulk Li (a), Na (b), Al (c), K (d), and Ca (e). The selected $\mathbf{q}$}-paths correspond to $H\Gamma N$ and $X\Gamma N$ for the bcc and fcc Bravais lattices, respectively. Such simple elemental metals show a parabolic-like dispersion of the main plasmon, with an effective mass that increases with the plasmon energy. Secondary excitations are more visible for heavier elements, showing a much flatter dispersion than that of the main plasmon.
Figure 5: Spectral band structures corresponding to the RPA loss function, $|\mathrm{Im}[Y(\omega,{\mathbf{q}})]|$, of cubic elemental metals from rows II-IV of the periodic table. The plots are made for two high-symmetry $\mathbf{q}$}-lines of bulk Li (a), Na (b), Al (c), K (d), and Ca (e).
Figure 6: Spectral band structures corresponding to the RPA loss function, $|\mathrm{Im}[Y(\omega,{\mathbf{q}})]|$, of cubic elemental metals from rows IV of the periodic table. The plots are made for two high-symmetry $\mathbf{q}$}-lines of bulk V (a), Cr (b), Fe (c), Ni (d), and Cu (e). A zoomed region is included in the lower panels.
Figure 7: Spectral band structures corresponding to the RPA loss function, $|\mathrm{Im}[Y(\omega,{\mathbf{q}})]|$, of hexagonal elemental metals from rows II-IV of the periodic table. The plots are made for four high-symmetry $\mathbf{q}$}-lines of bulk Be (a), Ti (b), and Co (c).
Fig. 6 shows analogous plots for V (a), Cr (b), Fe (c), Ni (d), and Cu (e), while Fig. 7 shows plots along the $HA\Gamma M|K\Gamma$ path for metals with hexagonal symmetry: Be (a), Ti (b), and Co (c). The complexity of the excitation landscape of the heavier metals increases very quickly with respect to the cubic elements in Fig. 5 and the light hexagonal ones (Be and Mg). The main plasmon is split into a multiple-peak structure with mostly flat dispersions, at variance with what is expected from the free electron gas model with an equivalent electronic density. There are also many more secondary excitations with a significant spectral weight, as discussed in the previous section. The dispersion of such excitations is also more complex, and gives rise to diverse features in the spectral band structure, as discussed in the following.
The plasmonic dispersions are in general anisotropic and may have discontinuities due to anisotropies in the electronic band structure [64, 82, 67, 83]. Examples of such discontinuities are visible in Fig. 7 at the $\Gamma$ point of the $A\Gamma M$ path, for instance, around $30$ eV for Be, 4545-$55$ eV for Ti, and 2020-$35$ eV for Co. In the latest case, the discontinuity extends to the region of the plasmon, similar to other hexagonal systems with $d$-electrons showing strong anisotropy signatures, such as ZnO [55]. Other spectral features are present in many of the studied metals, such as non-parabolic energy and intensity dispersions [84, 83], including negative dispersions and indirect excitations, i.e., with a finite intensity at finite $\mathbf{q}$}, that vanishes at ${\mathbf{q}}=\Gamma$. There is also quasi-particle overlapping leading to band crossings and anti-crossings, as shown in Fig. 6 in the zoomed region around $10$ eV.
Individual plots analogous to the ones in Figs. 5, 6, and 7, for all the 25 metals studied here, can be found in Ref. [85]. An effective analytical representation of the spectral $Y({\mathbf{q}},\omega)$ band structure and the interpretation of its features is presented in the next section.
III.5 Spectral properties within MPA($\mathbf{q}$})
Figure 8: Comparison of the spectral $Y({\mathbf{q}},\omega)$ band structure of V in the $\Gamma N$ direction, generated from the numerical RPA data with its corresponding MPA model with $n_{Y}=13$ poles.
Many of the studied metals exhibit spectral $Y({\mathbf{q}},\omega)$ band structures with a high degree of complexity, as illustrated in the previous section. Still, one can track the dispersion of the most prominent peaks, which varies smoothly with $\mathbf{q}$}. This suggests that such a complex spectrum can be modeled with a small set of plasmonic excitations with a simple $\mathbf{q}$} dependence.
To test this hypothesis, we have fitted MPA($\mathbf{q}$}) models, with a number of poles of $n_{Y}=2$ to $n_{Y}=15$, to the numerical $Y({\mathbf{q}},\omega)$ data of the 25 metals studied in this work. The fit is performed along a selected high symmetry $\mathbf{q}$}-line, $\Gamma N$ for cubic systems and $\Gamma M$ for hexagonal. Analogous fits can be made along the other $\mathbf{q}$}-paths. The results can be found in Ref. [85]. In all the cases, the resulting $Y^{\mathrm{MPA}}({\mathbf{q}},\omega)$ reproduces very accurately the numerical data, even for systems with complex pole structures.
Figure 9: Spectral $Y({\mathbf{q}},\omega)$ band structure of Ca reconstructed with MPA($\mathbf{q}$}) in Eqs. (14) and (15) with a number of $n_{Y}=6$ poles. The total spectral function along the $\Gamma N$ $\mathbf{q}$}-path is shown in (a), while the individual contribution of each pole is plotted in (b-g) according to their energy position at $\Gamma$. The dashed lines in (a) indicate the energy dispersion of $\mathrm{Re}[\Omega_{p}]$.
Fig. 8 shows the case of V, which is fitted with the same number of $n_{Y}=13$ poles than in Fig. 3. The numerical spectral band structure in Fig. 8(a) is accurately reproduced with MPA($\mathbf{q}$}) in Fig. 8(b), while Fig. 8(c) includes dashed black lines indicating the energy position of the poles, $\mathrm{Re}[\Omega_{p}({\mathbf{q}})]$. Analogous plots for the rest of the 25 metals are provided in Ref. [85]. The analytical MPA($\mathbf{q}$}) representation also offers the possibility to disentangle overlapping poles, which in many cases is not trivial. In the following, we use Ca and Ni to exemplify such untangling.
Fig. 9(a) shows the total spectral band structure of Ca, reconstructed with MPA($\mathbf{q}$}) with 6 poles, analogous to the case of V in Fig. 9(c). In panels (b-g) we show the contribution of each individual pole to the sum in Eq. (14). The poles are labeled according to their energy position at ${\mathbf{q}}=0$. The resulting MPA($\mathbf{q}$}) poles of Ca present a sufficiently small broadening and smoothly varying intensity to allow identifying them as well defined bands along the whole $\mathbf{q}$} range. The spectrum shows two main poles in line with Ref. [57]. For most values of $\mathbf{q}$}, the most intense one is $p=5$, while $p=4$ dominates around ${\mathbf{q}}=N$. Both poles present a non-monotonic spectral weight with a maximum at finite $\mathbf{q}$}. The three upper poles ($p=4,5,6$) cross at a $\mathbf{q}$}-point slightly before half the $\Gamma N$ distance. The maximum total intensity is located around this region, as a consequence of a constructive pole superposition: $\mathrm{Re}[R_{p}({\mathbf{q}})]>-\mathrm{Im}[R_{p}({\mathbf{q}})]>0$, reflecting the constraints in Eq. (16) for the pole positions $\Omega_{p}$.
Figure 10: Spectral $Y({\mathbf{q}},\omega)$ band structure of Ni reconstructed with MPA($\mathbf{q}$}) in Eqs. (14) and (15) with a number of $n_{Y}=13$ poles. The total spectral function along the $\Gamma N$ $\mathbf{q}$}-path is shown in (a), while the individual contribution of each pole is plotted in (b-n) according to their energy position at $\Gamma$. The dashed lines in (a) indicate the energy dispersion of $\mathrm{Re}[\Omega_{p}]$.
In Fig. 10, we show analogous MPA($\mathbf{q}$}) spectral band structures for Ni, in this case modeled with 13 poles. As for Ca, the most intense poles are well defined and superimpose constructively with the intensity spreading in a larger energy range. However, many poles, especially in the low-energy region (p=2p=2-66) around $10$ eV shown in Fig. 6, exhibit residues with an imaginary part larger than its real part and a sign not always complying with the time ordering. Imposing such constraints on the residues $R_{p}$, using a relation analogous to the one for the poles in Eq. (16), results in a reduction of the accuracy of the MPA($\mathbf{q}$}) fit with respect to the numerical data. The need for a fit with free residues, which allows for incoherent pole interference, evidences the complexity of plasmonic quasi-particle formation in non-homogeneous electronic materials.
IV Conclusions
We have presented a thorough study of the frequency and momentum dependent inverse dielectric function of 25 elemental metals, computed from first principles at the RPA level of the many-body theory. Such an approach reproduces well the experimental measurements. The RPA data are used to construct spectral band structures of plasmonic excitations. Such excitations exhibit a complex landscape in inhomogeneous metals, on the basis of the mixed orbital character of the valence states. The plasmonic dispersions are much flatter than expected from the free-electron gas model with an equivalent electronic density or plasmon energy, particularly when most of the spectral weight is distributed among several poles. The energy position of the main MPA pole is used to compute an effective number of electrons and compare with free-electron gas model. The spectral weight of such pole is used as a quantitative measure of the deviation from a single pole PPA dielectric function and provides a general picture of its variability across elemental metals.
We have also introduced MPA($\mathbf{q}$}), a generalized analytical representation of the dynamical part of the inverse dielectric function, as an effective multipole-Padé model depending on both momentum and frequency. With our generalized approach, we have accurately fitted the numerical data, with a limited number of poles ranging from 2 to 15 for all the systems. The obtained MPA($\mathbf{q}$}) representations describe complex features present in the spectral band structures, such as non-parabolic energy and intensity dispersions and constructive and destructive quasiparticle overlapping, leading to band crossings and anti-crossings. We report the specific MPA($\mathbf{q}$}) model for each elemental metal, which can be used as a building block for other works, as a first-principle alternative to over simplified analytical models such as the Drude/Lindhard dielectric function. Moreover, the MPA($\mathbf{q}$}) representation in both momentum and frequency can be fundamental to reduce the computational cost of GWGW/BSE calculations, as a starting point towards GWGW/BSE performed on a $\mathbf{k}$$\mathbf{q}$}-path, which would considerably reduce the computational cost of such calculations.
Acknowledgments
The computations of this work were carried out on UNINETT Sigma2 high performance computing resources (grant NN9650K), and through the project EHPC-EXT-2022E01-022 Joint Undertaking, that granted access to Leonardo-Booster@Cineca, Italy, provided by EuroHPC. Work in Norway is supported by the Research Council of Norway in the MORTY project (664350).
Data Availability
The data generated in this article are openly available.